OpenStudy (anonymous):

1. A line passes through the point (–6, 1) and has slope m = –4. Find the equation in point–slope form.

OpenStudy (anonymous):

point slope form: y-y1=m(x-x1) (x1,y1) is the point and m is the slope. just plug those in

OpenStudy (anonymous):

So it's B. y - 4 = -1(x + 6) @CalebBeavers

OpenStudy (anonymous):

No, y1 is the y coordinate of your point, x1 is the x coordinate and m is your slope. y-1=-4(x+6)

OpenStudy (anonymous):

y - 1 = -4(x + 6) @CalebBeavers